OpenStudy (anonymous):

Use complete sentences to explain how you would use graphing technology to find the solutions of 0 = 2x2 − 3x + 0.5. What do the solutions of 0 = 2x2 − 3x + 0.5 represent on the graph of the equation?

OpenStudy (anonymous):

The problem you listed is a quadratic. This means it looks like a big U, either upside down or right side up. The thing about quadratics is, if you graph it. It has one point that is either and absolute max or min. That is the point where the graph changes slope from pos/neg or vice versa. The x values are the point where the graph intercepts the x-axis when y = 0. Remember there will be two points. So if we drew this out, you could find it intercepts the x axis twice. Once at x = 0.014 and again at x = 17.486.

OpenStudy (anonymous):

Here is the correct graph for your problem

OpenStudy (anonymous):

Normally the equation is written as y = 2x^2 - 3x + 0.5. As you can see, you set y = 0. You are essentially saying to find all the x values when y =0. Had you put it a different value of y, you would have different x values. I know you didn't ask for this, but if you want to find the minimum or max (this graph has a min) you would use the equation x = -b/2a. Remember the formula ax^2 + bx + c? So for your equation a - 2, b = -3. That gives you x= 3/2. That point is the middle of your graph. Plug that in for x, and you can find your minimum too. Hope that helped.